# Split & Merge: Unlocking the Potential of Visual Adapters via Sparse Training

### 05 Dec 2023 v1

@@ 4.4 Visualization
Figure 5: Attention map visualization.
Here, we use attention maps to visually illustrate the importance of the hierarchical sparse strategy. It can be observed that compared to AdaptFormer, our MoSA exhibits more focused attention on foreground objects (indicated by darker colors in Fig. 5), and the hierarchical sparse strategy further strengthens this attention. In the hierarchical sparse strategy, the dense down-projection layer provides robust intermediate features by receiving the entire training data, while multiple sparse up-projection layers project features into different subspaces for different data. The combination of the two components allows for better attention.
@@ 5 Conclusion
In this paper, we focus on the Adapter Tuning method in parameter-efficient fine-tuning and propose MoSA to improve the performance of standard adapters without any extra parameters or computation. Recognizing that standard adapters still suffer from parameter redundancy, we combine sparse training with stochastically activated MoE to fully utilize all parameters within the adapters. Comprehensive experiments on a total of 27 datasets show that MoSA consistently outperforms all other baselines, achieving state-of-the-art performance in Adapter Tuning. We hope that our work could inspire researchers to reconsider the issue of parameter redundancy in adapters and make further advancements towards more efficient PEFT methods.
Supplementary Material
In this supplementary material, we present detailed information about the datasets and implementation of the experiments, along with additional experimental results of MoSA across different structures and additional visualizations. In Sec. A, we show details about all the datasets used in experiments. In Sec. B, we provide details about the experimental configurations. In Sec. C, we conduct additional experiments of MoSA on different backbone scales, model architectures, adapter structures, and PEFT methods. In Sec. D, we give additional visualization results.
@@ A Dataset details
Table 8: Details of all datasets used to validate MoSA.
Here, we describe the details of all datasets used to validate MoSA. The number of classes and the train/valid/test splits for each dataset are shown in Tab. 8.
B Implementation details
@@
Table 9: Implementation details for Adapter Tuning and LoRA.
In Tab. 9, we summarize all experimental configurations with Adapter Tuning and LoRA. For other baselines, we just follow the settings in [29]. Following the linear scaling rule  [9, 17, 35, 23], the learning rate is set as $base\_lr\times b/256$, where $b$ is the batch size and $base\_lr$ is chosen from the range specified in Tab. 9.
C More experiments
In this section, we validate the performance of MoSA on different backbone scales, model architectures, adapter structures, and PEFT methods. We also show the pre-task results for MoSA with ViT-B/16 on VTAB-1k. Similar to the main text, Top-1 accuracy (%) is used to evaluate the performance of the methods on the respective datasets, and the number (M) of extra parameters (trainable parts excluding the classifier) is used to assess the efficiency of the methods. The best accuracy is highlighted in bold, while the second one is underlined.
Full fine-tuning 88.3 85.9 96.7 93.1 86.8 90.16 303.30
Linear probing 84.7 78.9 97.4 89.6 55.1 81.14 0.00
BitFit [74] 88.5 86.4 98.8 93.5 83.2 90.08 0.27
AdaptFormer-64 [6] 89.3 86.3 98.8 93.2 80.9 89.70 3.17
SparseAdapter-64 [24] 89.4 86.8 99.0 93.9 82.3 90.28 0.79
MoSA-64 (Ours) 89.7 87.2 99.4 94.5 84.9 91.06 3.17
Table 10: Results on FGVC with ViT-L/16 backbone pre-trained on ImageNet-21K.
MoSA on different backbone scales. Here we evaluate MoSA with a larger backbone ViT-L/16 (303.3M vs. 85.8M ViT-B/16). The results in Tab. 10 indicate that, with the increase in backbone scale, the performance of the standard adapter cannot surpass full fine-tuning (89.70% vs. 90.16%). Sparse training could improve the performance of adapters , leading by 0.58% and 0.12% over the standard adapter and full fine-tuning, respectively. Our MoSA further enhances the performance by 0.78%, outperforming full fine-tuning by a large margin.
@@ C More experiments
MoSA on different model architectures. In addition to the standard ViT, we also experiment with another hierarchical vision Transformer, Swin-B [43], to demonstrate the effectiveness of MoSA. Similar to ViT, we can easily apply MoSA to Swin. As shown in Tab. 11, due to the strong feature extraction capability of this model architecture, full fine-tuning performs well on Swin, while other PEFT methods show suboptimal performance. It’s worth noting that our MoSA is the only method that outperforms full fine-tuning on Swin (90.72% vs. 90.14%), indicating that MoSA consistently adapts various vision Transformers to downstream tasks and improves performance.
@@
Table 12: Results on FGVC with different adapter structures.
MoSA on different adapter structures. As a supplement, we apply MoSA to two different adapter structures: Pfeifferr [57] and Houlsby [26]. Both Pfeifferr and Houlsby use a sequential connection for adapter design, with Pfeifferr incorporating the adapters only after the FFN layers, while Houlsby includes the adapters after both the Attention and FFN layers. The performance comparison on the two adapter structures is shown in Tab. 12. In this experiment, the bottleneck dimension for all adapters is set to 64. It can be observed that on both structures, sparse training brings improvements of 1.50% and 1.30% over the standard adapter, and the corresponding versions of our MoSA further yield performance gains of 1.70% and 1.76%.
LoRA-16 [27] 87.2 83.5 98.6 89.3 83.7 88.44 0.59
SparseLoRA-16 [24] 87.4 84.9 98.9 91.1 79.9 88.44 0.15
MoSL-16 (Ours) 89.0 85.6 99.3 91.8 83.9 89.92 0.59
Table 13: Results on FGVC with LoRA.
MoSA on different PEFT methods. To demonstrate the generality of our approach, we choose another widely used PEFT method, namely LoRA [27]. LoRA achieves parameter-efficient fine-tuning by applying a low-rank decomposition to the weight updates of linear layers. During inference, the additional modules could be merged into the pre-trained model parameters, resulting in zero extra overhead during inference. In Tab. 13, we compare the LoRA version sparse adapter (namely SparseLoRA) and our MoSA (namely MoSL) with the standard LoRA, with the rank of all LoRA modules set to 16. Our MoSL outperforms both the standard and sparse versions of LoRA, achieving an improvement of 1.48%, further confirming the effectiveness and rationality of our design.
Method Dataset Natural Specialized Structured VTAB
@@
Mean Acc. (%)
Mean Params. (M)
Full fine-tuning 68.9 87.7 64.3 97.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 65.57 85.80
Linear probing 63.4 85.0 64.3 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 53.00 0.00
BitFit [74] 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 62.05 0.10
AdaptFormer [6] 78.9 90.0 67.0 98.7 89.0 72.2 53.2 81.5 95.7 81.2 75.2 70.6 57.4 39.3 70.6 54.5 42.4 25.3 33.3 67.16 0.30
SparseAdapter [24] 79.1 89.2 65.7 98.6 89.3 68.5 52.5 79.5 94.7 79.4 74.4 70.2 56.9 37.8 70.9 51.3 41.3 25.2 32.5 66.16 0.08
MoSA (Ours) 79.7 91.5 66.2 98.8 89.7 79.0 53.4 83.4 95.6 82.0 75.1 71.5 58.1 40.7 70.2 57.8 43.6 26.5 34.0 68.25 0.30
Table 14: Performance comparisons on VTAB-1k with ViT-B/16 models pre-trained on ImageNet-21K.
Per-task results for MoSA on VTAB-1k. Tab. 14 shows the per-task results of MoSA on VTAB-1k. It can be seen that MoSA outperforms full fine-tuning on 13 tasks of VTAB, the highest among all PEFT methods. Additionally, MoSA also surpasses full fine-tuning (68.25% vs. 65.57%) and other baselines (AdaptFormer 67.16%) in the average accuracy across 19 tasks.
D More visualizations
Figure 6: Attention map visualization on FGVC. From left to right: the RGB image, AdaptFormer, our MoSA and Hierarchical MoSA.
Attention map. Here, we present more attention map visualizations in Fig. 6. Compared to AdaptFormer, our MoSA focuses more on foreground objects and exhibits more concentrated attention (highlighted in red in the images). And the hierarchical sparse strategy further strengthens this attention, providing intuitive evidence of the effectiveness of our design.
Figure 7: t-SNE visualization on CIFAR-100. From left to right: full fine-tuning, linear probing, BitFit, AdaptFormer, SparseAdapter and our MoSA (best viewed in color).
t-SNE [64]. We also provide t-SNE visualizations to show the feature distribution of different methods on the CIFAR-100 dataset. The results are shown in Fig. 7. We can observe that our MoSA achieves better feature clustering compared to full fine-tuning and other PEFT baselines.
E Pipeline
Algorithm 1 Pipeline of MoSA
0:   adapter paramters $w$, expert number $N$
   $\{w_{i}\}_{i=1}^{N}\leftarrow\text{Split}(w)$;
   for each epoch do
      for each training batch $x$ do
         $w_{j}\leftarrow\text{StochasticActive}(\{w_{i}\}_{i=1}^{N})$;
         $\text{ForwardOneBatch}(w_{j},x)$;
      end for
      $w^{*}\leftarrow\text{Merge}(\{w_{i}\}_{i=1}^{N})$;
      $\text{Evaluate}(w^{*},x^{\text{test}},y^{\text{test}})$;
   end for
We demonstrate the pipeline of our MoSA with pseudo-code in Algorithm 1.

### 23 Mar 2024 v2

@@ 4.5 Visualization
Here, we provide t-SNE visualizations to show the feature distribution of different methods on the CIFAR-100 dataset. The results are shown in Figure 4. We can observe that our MoSA achieves better feature clustering compared to full fine-tuning and other PEFT baselines.
@@ 5 Conclusion
In this paper, we focus on the Adapter Tuning method in parameter-efficient fine-tuning and propose MoSA to improve the performance of standard adapters without any extra parameters or computation. Recognizing that standard adapters still suffer from parameter redundancy, we combine sparse training with multiple stochastically activated experts to fully utilize all parameters within the adapters. Comprehensive experiments on a total of 27 datasets show that MoSA consistently outperforms all other baselines, achieving state-of-the-art performance in Adapter Tuning. We hope that our work could inspire researchers to reconsider the issue of parameter redundancy in adapters and make further advancements towards more efficient PEFT methods.
@@ A Dataset Details
Here, we describe the details of all datasets used to validate MoSA. The number of classes and the train/valid/test splits for each dataset are shown in Table 11.
B Implementation Details
In Table 12, we summarize all experimental configurations with Adapter Tuning and LoRA. For other baselines, we just follow [36]. Following the linear scaling rule  [10, 22, 42, 28], the learning rate is set as $base\_lr\times b/256$, where $b$ is the batch size and $base\_lr$ is chosen from the range specified in Table 12.
Table 11: Details of all datasets used to validate MoSA.
@@ B Implementation Details
Table 12: Implementation details for Adapter Tuning and LoRA.
@@
C More Experimental Results
In this section, we validate MoSA on different backbone architectures and adapter structures. We also show the pre-task results for MoSA with ViT-B/16 on VTAB-1k. Similar to the main text, Top-1 accuracy (%) is used to evaluate the performance of the methods on the respective datasets, and the number (M) of extra parameters (trainable parts excluding the classifier) is used to assess the efficiency. The best accuracy is highlighted in bold, while the second one is underlined. The results of our method is highlighted with a red background.
@@ C More Experimental Results
MoSA on different model architectures. In addition to the standard ViT, we also experiment with another hierarchical vision Transformer, Swin-B [51], to demonstrate the effectiveness of MoSA. Similar to ViT, we can easily apply MoSA to Swin. As shown in Table 13, due to the strong feature extraction capability of this model architecture, full fine-tuning performs well on Swin, while other PEFT methods show suboptimal performance. It’s worth noting that our MoSA is the only method that outperforms full fine-tuning on Swin (90.72% vs. 90.14%), indicating that MoSA consistently adapts various vision Transformers to downstream tasks and improves performance.
Table 14: Results on FGVC with different adapter structures.
@@
MoSA on different adapter structures. As a supplement, we apply MoSA to two different adapter structures: Pfeifferr [64] and Houlsby [33]. Both Pfeifferr and Houlsby use a sequential connection for adapter design, with Pfeifferr incorporating the adapters only after the FFN layers, while Houlsby includes the adapters after both the Attention and FFN layers. The performance comparison on the two adapter structures is shown in Table 14. In this experiment, the bottleneck dimension for all adapters is set to 64. It can be observed that on both structures, sparse training brings improvements of 1.50% and 1.30% over the standard adapter, and the corresponding versions of our MoSA further yield performance gains of 1.70% and 1.76%.
Table 15: Per-task results on VTAB-1k with ViT-B/16 pre-trained on ImageNet-21K.
Method Dataset Natural Specialized Structured
@@
Mean
Full fine-tuning 68.9 87.7 64.3 97.2 86.9 87.4 38.8 75.88 79.7 95.7 84.2 73.9 83.36 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 65.57
Linear probing 63.4 85.0 64.3 97.0 86.3 36.6 51.0 68.93 78.5 87.5 68.6 74.0 77.16 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 53.00
BitFit [80] 72.8 87.0 59.2 97.5 85.3 59.9 51.4 73.30 78.7 91.6 72.9 69.8 78.25 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 62.05
VPT-shallow [36] 77.7 86.9 62.6 97.5 87.3 74.5 51.2 76.81 78.2 92.0 75.6 72.9 79.66 50.5 58.6 40.5 67.1 68.7 36.1 20.2 34.1 64.85
VPT-deep [36] 78.8 90.8 65.8 98.0 88.3 78.1 49.6 78.48 81.8 96.1 83.4 68.4 82.43 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8 69.42
AdaptFormer [7] 78.9 90.0 67.0 98.7 89.0 72.2 53.2 78.42 81.5 95.7 81.2 75.2 83.41 70.6 57.4 39.3 70.6 54.5 42.4 25.3 33.3 67.16
SparseAdapter [30] 79.1 89.2 65.7 98.6 89.3 68.5 52.5 77.58 79.5 94.7 79.4 74.4 81.99 70.2 56.9 37.8 70.9 51.3 41.3 25.2 32.5 66.16
MoSA (Ours) 79.7 91.5 66.2 98.8 89.7 79.0 53.4 79.86 83.4 95.6 82.0 75.1 84.03 71.5 58.1 40.7 70.2 57.8 43.6 26.5 34.0 68.25
Per-task results for MoSA on VTAB-1k. Table 15 shows the per-task results of MoSA on VTAB-1k. It can be seen that MoSA outperforms full fine-tuning on 13 tasks of VTAB, the highest among all PEFT methods. Additionally, MoSA also surpasses full fine-tuning (68.25% vs. 65.57%) and other baselines (AdaptFormer 67.16%) in the average accuracy across 19 tasks.